OpenStudy (anonymous):

Evaluate the function f(x) = x 2 + 1 for f(2). a.1 b.5 c.7

OpenStudy (agent0smith):

f(2) means plug in 2 for x, and evaluate

OpenStudy (anonymous):

what would be the answer?

OpenStudy (agent0smith):

lol, the answer will be what you work out when you plug in x=2. since f(x) = x^2 + 1, then... f(2) = 2^2 + 1

OpenStudy (anonymous):

so it 5 right can you help me out with a few more?

OpenStudy (agent0smith):

5 is correct. Okay.

OpenStudy (anonymous):

can you guys help me with one more its If g(x) = 3x + 1, evaluate g(-2). a.-5 b.2 c.5